OpenStudy (anonymous):

3. Simplify 3n(4n^2 + 5n)

OpenStudy (diyadiya):

Multiply each term \(3n(4n^2 + 5n) \) \(3n*4n^2 + 3n*5n\) What do you get now?

OpenStudy (anonymous):

can you give the answer please

OpenStudy (diyadiya):

No i cant give the Final answer you must try :)

OpenStudy (anonymous):

7n^2+8n

OpenStudy (diyadiya):

You are multiplying 3n with 4n^2 \[3n \times 4n^2= 3\times 4 \times n \times n^2 =12 n^3\] Note :\[\huge n^a \times n^b=n^{a+b}\] Now can you do 3n*5n

OpenStudy (anonymous):

12n^2 + 8n

OpenStudy (diyadiya):

What is \( 3n \times 5n\) ? \( 3 \times 5 \times n \times n \) Now can you multiply and tell me What you get?

OpenStudy (anonymous):

15n^2

OpenStudy (diyadiya):

Right ! Now what's the final answer?

OpenStudy (anonymous):

12n^3+15n^2

OpenStudy (diyadiya):

Yeah ! Good

OpenStudy (anonymous):

thank you